What is 1% of 77.2
Answer by fcabanski(1391)

1% = 1/100 = .01.


x% of something = x/100 * something.


1% of 77.2 = 1/100 * 77.2 or .01 * 77.2 = .772

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com